# Problem-Solving in Language Model Networks

### v1

@@ Abstract
To improve the reasoning and question-answering capabilities of Large Language Models (LLMs), several multi-agent approaches have been introduced. While these methods enhance performance, the application of collective intelligence-based approaches to complex network structures and the dynamics of agent interactions remain underexplored. This work extends the concept of multi-agent debate to more general network topologies, measuring the question-answering accuracy, influence, consensus, and the effects of bias on the collective. The results show that random networks perform similarly to fully connected networks despite using significantly fewer tokens. Furthermore, a strong consensus among agents in correlates with correct answers, whereas divided responses typically indicate incorrect answers. Analysing the influence of the agents reveals a balance between self-reflection and interconnectedness; self-reflection aids when local interactions are incorrect, and local interactions aid when the agent itself is incorrect. Additionally, bias plays a strong role in system performance with correctly biased hub nodes boosting performance. These insights suggest that using random networks or scale-free networks with knowledgeable agents placed in central positions can enhance the overall performance of multi-agent systems.
@@ Methods
(b) Follow-up prompt for the agent to reconsider their answer after being exposed to their own and others’ responses. This prompt is concatenated with (a) for QA in follow-up rounds. Here self_response refers to the previous response of the individual agent and neighbour_response refers to the concatenation of responses of the agent’s neighbours.
@@
Figure 3: The prompt used to generate reasoning for biased agents, where question is the multiple-choice question and biased_answer is either the correct answer for correctly biased agents or an incorrect answer for incorrectly biased agents.
@@ Results
To understand how an agent may be influenced, the probability of the agent being correct in round $n$, given its previous response and the response of its neighbours in round $n-1$ is shown in Fig 7, with green (red) points showing cases where the agents own response was correct (incorrect) in the previous round. For clarity, the figure depicts the case of random networks, where each node will have, on average, the same number of neighbours. The figure shows that as the number of correct neighbours increases, so too does the probability of the agent being correct. Furthermore, the tendency for green points to lie above red points highlights the positive impact of self-reflection; regardless of the neighbour’s response, an agent is more likely to answer correctly if it was correct in the previous round.
Considering the edge cases reveals several interesting insights about the agent’s behaviour. Firstly, when the agent and all of its neighbours were incorrect in the previous round, there is a $10\%$ chance that the agent will spontaneously switch to the correct answer.‘ If the agent was correct on the other hand, this probability increases to $30\%$ demonstrating the positive effects of self-reflection. Secondly, when an agent was incorrect in the previous round, but is surrounded by correct agents, it is very unlikely to remain incorrect, showing a decrease in the impact of self-reflection. Finally, in the case where the neighbour’s responses are split, the agent is equally likely to be correct or incorrect, with a slight increase for agents correct in the previous round. These findings highlight the importance of both individuality and collective thinking in multi-agent systems. That is, collaborative problem-solving improves the overall performance of the collective, while self-reflection acts to improve performance when local interactions are misguided.
@@
Considering bias, networks correctly biased at their hubs (Fig 8(e)) exhibit a large number of agents switching from incorrect to correct answers after the first round, in agreement with Fig 6. These agents with correct answers tend to keep the correct response throughout the remaining rounds of debate. Networks incorrectly biased at their hubs (Fig 8(f)), on the other hand, have an increasing number of agents switching from correct to incorrect after round two. This is a significant result, as it highlights the fact that agents may have the correct answer, but will be convinced to switch due to the influence of their biased neighbours. Moreover, this suggests that multi-agent influence plays a larger role in performance than self-reflection, in general agreement with Fig 7. If self-reflection had a significant impact on the system then correct agents would disregard the responses of their incorrect neighbours, highlighting the importance of collective intelligence and collaboration in the problem-solving domain. As expected, networks biased at their edges (Fig 8(g) and Fig 8(h)) have similar dynamics to the unbiased counterpart.

### v2

@@ Abstract
To improve the reasoning and question-answering capabilities of Large Language Models (LLMs), several multi-agent approaches have been introduced. While these methods enhance performance, the application of collective intelligence-based approaches to complex network structures and the dynamics of agent interactions remain underexplored. This work extends the concept of multi-agent debate to more general network topologies, measuring the question-answering accuracy, influence, consensus, and the effects of bias on the collective. The results show that random networks perform similarly to fully connected networks despite using significantly fewer tokens. Furthermore, a strong consensus among agents correlates with correct answers, whereas divided responses typically indicate incorrect answers. Analysing the influence of the agents reveals a balance between self-reflection and interconnectedness; self-reflection aids when local interactions are incorrect, and local interactions aid when the agent itself is incorrect. Additionally, bias plays a strong role in system performance with correctly biased hub nodes boosting performance. These insights suggest that using random networks or scale-free networks with knowledgeable agents placed in central positions can enhance the overall question-answering performance of multi-agent systems.
@@ Methods
(b) Follow-up prompt for the agent to reconsider their answer after being exposed to their own and others’ responses. This prompt is concatenated with (a) for QA in follow-up rounds. Here self_response refers to the previous response of the individual and neighbour_responses refers to the concatenation of the neighbours’ responses.
@@
Figure 3: The prompt used to generate reasoning for biased agents, where biased_answer is either the correct answer for correctly biased agents or an incorrect answer for incorrectly biased agents.
@@ Results
To understand how an agent may be influenced, the probability of the agent being correct in round $n$, given its previous response and the response of its neighbours in round $n-1$ is shown in Fig 7, with green (red) points showing cases where the agents own response was correct (incorrect) in the previous round. For clarity, the figure depicts the case of random networks, where each node will have, on average, the same number of neighbours. The figure shows that as the number of correct neighbours increases, so too does the probability of the agent being correct. Furthermore, the tendency for green points to lie above red points highlights the positive impact of self-reflection; regardless of the neighbours’ responses, an agent is more likely to answer correctly if it was correct in the previous round.
Considering the edge cases reveals several interesting insights about the agent’s behaviour. Firstly, when the agent and all of its neighbours were incorrect in the previous round, there is a $10\%$ chance that the agent will spontaneously switch to the correct answer. If the agent was correct on the other hand, this probability increases to $30\%$ demonstrating the positive effects of self-reflection. Secondly, when an agent was incorrect in the previous round, but is surrounded by correct agents, it is very unlikely to remain incorrect, showing a decrease in the impact of self-reflection. Finally, in the case where the neighbours’ responses are split, the agent is equally likely to be correct or incorrect, with a slight increase for agents correct in the previous round. These findings highlight the importance of both individuality and collective thinking in multi-agent systems. That is, collaborative problem-solving improves the overall performance of the collective, while self-reflection acts to improve performance when local interactions are misguided.
@@
Considering bias, networks correctly biased at their hubs (Fig 8(e)) exhibit a large number of agents switching from incorrect to correct answers after the first round, in agreement with Fig 6. These agents with correct answers tend to keep the correct response throughout the remaining rounds of debate. Networks incorrectly biased at their hubs (Fig 8(f)), on the other hand, have an increasing number of agents switching from correct to incorrect after round two. This is a significant result, as it highlights the fact that agents may have the correct answer, but will be convinced to switch due to the influence of their biased neighbours. Moreover, this suggests that multi-agent influence plays a larger role in performance than self-reflection, in general agreement with Fig 7. If self-reflection had a significant impact on the system then correct agents would disregard the responses of their incorrect neighbours, highlighting the importance of collaboration in the problem-solving domain. As expected, networks biased at their edges (Fig 8(g) and Fig 8(h)) have similar dynamics to the unbiased counterpart.